A job order cost sheet for Ryan Company is shown below.A job order cost sheet for Ryan Company is shown b
Answer #1

a.1) Balance in Work in Process Inventory on January 1 = Direct Materials + Direct labor + Manufacturing overhead

= 6,600 + 6,700 + 4,690

= 17,990

2) Last year overhead rate = Manufacturing overhead / Direct labor

= 4,690 / 6,700

= 70%

Current year overhead rate = Manufacturing overhead / Direct labor

= 8,000 / 10,000 = 80%

or

= 3,840 / 4,800 = 80%

b)

Date Account titles and explanation Debit Credit
Jan.31 Work in process inventory 9,000
Raw materials inventory 9,000
(To record raw materials used)
31 Work in process inventory 14,800
Factory labor 14,800
(To record factory labor used)
31 Work in process inventory 11,840
Manufacturing overhead 11,840
(To record manufacturing overhead)
31 Finished Goods Inventory 53,630
Work in Process Inventory 53,630
(To record job completed)